OpenStudy (anonymous):

Which of the following is the simplified form of x plus 4 over 2 x plus 6 plus the fraction 3 over x squared minus 9 ?

OpenStudy (anonymous):

\[\frac{ x - 4 }{ 2x + 6 } + \frac{ 3 }{ ?x^2 - 9}\]

OpenStudy (whpalmer4):

\[\frac{x+4}{2x+6}+\frac{3}{x^2-9}\]?

OpenStudy (anonymous):

Oops, there's a question mark in there, and it shouldn't be there

OpenStudy (whpalmer4):

is x-4 or x+4?

OpenStudy (whpalmer4):

in any case, start by factoring everything as much as possible, then make a common denominator for the two fractions so you can combine them

OpenStudy (anonymous):

x-4

OpenStudy (whpalmer4):

okay, because your problem statement says "x plus 4"

OpenStudy (anonymous):

I'm sorry, it is x+4

OpenStudy (anonymous):

I got this far: \[\frac{ x+4 }{ 2(x+3) } + \frac{ 3 }{ (x-3)(x+3) }\]

OpenStudy (whpalmer4):

okay, to make the common denominator you'll have to multiply the lh fraction by (x-3)/(x-3) and the rh fraction by 2/2

OpenStudy (whpalmer4):

do you see why?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

good. have fun :-)

OpenStudy (anonymous):

is it: \[\frac{ 6+(x+4) }{ 2(x+3) }\]

OpenStudy (anonymous):

\[x + 2 \over 2(x + 3)\] \[ x - 2 \over 2(x - 3) \] \[ x - 2 \over 2(x +3)\] \[ x+ 2 \over 2(x- 3) \]

OpenStudy (whpalmer4):

Hmm, looks like you went off the rails somewhere... \[\frac{ x+4 }{ 2(x+3) } + \frac{ 3 }{ (x-3)(x+3) } = \frac{(x+4)} {2(x+3)}*\frac{(x-3)}{(x-3)} + \frac{3}{(x-3)(x+3)}*\frac{2}{2}\]\[=\frac{(x+4)(x-3) + 3*2}{2(x-3)(x+3) } = \frac{ x^2-3x+4x-12+6}{2(x-3)(x+3) }=\frac{x^2+x-6} {2(x-3)(x+3)}\]Now we factor the numerator, cancel common factors and we're done.

OpenStudy (anonymous):

Its adding not multiplication

OpenStudy (whpalmer4):

You have to multiply to get the common denominator in order to add

OpenStudy (whpalmer4):

just like adding 1/4 and 1/7

OpenStudy (whpalmer4):

\[\frac{1}{4} + \frac{1}{7} = \frac{1}{4}*\frac{7}{7} + \frac{1}{7}*\frac{4}{4} = \frac{7}{28} + \frac{4}{28} = \frac{11}{28}\]

OpenStudy (anonymous):

But it has to be simplified so the (x-3)s would cancel each other, leaving: 6+(x+4)/2(x+3)

OpenStudy (whpalmer4):

you're doing it wrong. what do you get when you factor \(x^2+x-6\)?